OpenStudy (harley1312):

What statement makes the open sentence 7x-4=5x+14 true? A. x=8 B. x=6 C. x=9 D. x=12

OpenStudy (xmissalycatx):

Hi there! Well first of all, what do you think it is?

OpenStudy (xmissalycatx):

And what you'd need to do is try to plug in each answer choice into where it says "x" in the equation..

OpenStudy (harley1312):

I don't even have an idea to start out with

OpenStudy (xmissalycatx):

Okay, like I said earlier, get each answer choice and plug them into the equation.. so for the first answer choice, if you plug it in you'd get 7 (8) - 4 = 5 (8) +14 ______________________ 56 - 4 = 40 + 14 Is that true?

OpenStudy (mattga2002):

Move all of the variables to one side and the integers to another. Simplify. If you do, you will find that the answer is the only one that makes sense

OpenStudy (mattga2002):

Step 1: Subtract 5x from both sides. 7x−4−5x=5x+14−5x 2x−4=14 Step 2: Add 4 to both sides. 2x−4+4=14+4 2x=18 Step 3: Divide both sides by 2. 2x 2 = 18 2

OpenStudy (mattga2002):

Sorry about step 3

OpenStudy (mattga2002):

\[2x/2 = 18/2\]

OpenStudy (mattga2002):

Solve for X

OpenStudy (xmissalycatx):

o.o lel

OpenStudy (jv0408):

has someone already answered your question?

OpenStudy (jv0408):

its going to be C